A pyrex glass flask is filled with 200 g of water. The energy required to heat the system (the flask and the water) from 20 °C to 45 °C is 23.1 kJ. Assume that the specific heat capacity of water is 4.18 J g–1 °C–1. What is the heat capacity of the flask?

  •  A. 36 J °C–1
  •  B. 88 J °C–1
  •  C. 76 J °C–1
  •  D. 60 J °C–1
  •  E. 2200 J °C–1
1
Expert's answer
2021-03-08T07:10:28-0500

mass of water, m=200g

H= 23.1Kj

T1= 20°C

T2= 45°C

C of water = 4.18J/g°C

But the mass of the glass was not given in the question, so let us assume that the combined mass of the glass and the water is the 200g given.

C of glass Cg= ?

H= m(T2-T1)(Cw + Cg)

23100= 200(45-20)(4.18+Cg)

23100= 5000(4.18+ Cg)

23100 = 5000Cg + 20900

23100-20900= 5000Cg

Cg= 2200/5000 = 0.44J/g°C


Unless the mass of the pyrex glass is given, there is no other alternative way of solving this question